40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17159     Entry time: Mon Sep 26 11:39:37 2022     In reply to: 17156     Reply to this: 17174
Author: Paco 
Type: Update 
Category: BHD 
Subject: BH55 RFPD installed - part II 

[Paco, Anchal]

We followed rana's suggestion for stress relief on the SMA joint in the BH55 RFPD that Radhika resoldered. We used a single core, pigtailed wire segment after cleaning up the solder joint on J7 (RF Out) and also soldered the SMA shield to the RF cage (see Attachment #1). This had a really good effect on the rigidity of the connection, so we moved back to the ITMY table.

We measured the TEST in to RF Out transfer function using the Agilent network analyzer, just to see the qualitative features (resonant gain at around 55 MHz and second harmonic suppression at around 110 MHz) shown in Attachment #2. We used 10kOhm series resistance in test input path to calibrate the measured transimpedance in V/A. The RFPD has been installed in the ITMY table and connected to the PD interface box and IQ demod boards in the LSC rack as before.

Measurement files

Attachment 1: PXL_20220926_175010061.MOTION.jpg  1.823 MB  Uploaded Mon Sep 26 12:49:38 2022  | Hide | Hide all
PXL_20220926_175010061.MOTION.jpg
Attachment 2: BH55_Transimpedance_Measurement.pdf  26 kB  Uploaded Mon Sep 26 13:08:56 2022  | Hide | Hide all
BH55_Transimpedance_Measurement.pdf
ELOG V3.1.3-